# Neural Networks for Extreme Quantile Regression with an Application to Forecasting of Flood Risk

### v3

@@ 1 Introduction
Figure 1 shows the river catchment of a gauging station on the Aare river in Bern, Switzerland. It is part of the Aare–Rhein basin, where flooding is a major economic and safety concern (Andres et al., 2021). The Swiss Federal Office for the Environment (FOEN) provides recordings of daily average discharges throughout the country. For risk assessment, they report the 100-year return levels using the GEV method for annual maxima and the GPD for large daily discharges. The horizontal lines in Figure 2 show estimates of this return level at the Bernese station based on data from the years 1930–1958. Both methods give very similar results.
@@
The disadvantage of such an unconditional approach is twofold. First, the return level is static and unable to reflect changes in the size of extreme floods over time, which can occur for instance due to climate change. For the Bernese station, for instance, a structural break has been observed in the nineties without a clearly defined cause111see flood report of the FOEN at https://www.hydrodaten.admin.ch/en/2135.html, making classical extreme value modelling challenging. Second, while the return level $Q^{T}$ is relevant for the construction of long-term flood infrastructure, it can not be used to assess the risk of flooding on a given day. Such forecasting of extreme events is crucial for early warning systems. Indeed, the probability of exceeding on a particular day a given high threshold, say the (constant) 100-year return level, depends on many covariates $\mathbf{X}$ such as the river flows upstream and precipitation in the catchments during the preceding days and weeks.
@@
The top panel of Figure 2 shows one-day-ahead forecasts of such conditional 100-year quantiles $Q^{100}_{\mathbf{x}}$ for the Bernese river data from the method of this paper, fitted to the years 1930–1958. We see that, as opposed to the unconditional return level $Q^{100}$, the conditional quantile changes from day to day depending on past precipitation and river flows. In fact, on August 21, the day before the first exceedance of the unconditional $Q^{100}$ during the serious flood of August 2005 in Bern, the size of the conditional 100-year event predicted for the next day was much higher than on other days, and the forecasted conditional probability of such an exceedance (bottom panel of Figure 2) was 920 times larger than the static 100-year probability. Both outputs can be used as triggers for early warnings and additional flood management measures. In particular, the days when an exceedance above $Q^{100}$ is likely can be effectively pinpointed thanks to their temporal sparsity in the probability forecast.
@@
Forecasting extreme events is notoriously difficult due to the low occurrence probabilities involved. This is particularly true for hydrological models, which are typically used by national agencies, and which do not use explicit tail extrapolation. In the aftermath of the 2005 flood, the FOEN published a detailed analysis of the internal forecasting procedures during this event (Bezzola and Hegg, 2007). The report showed that too-late warnings lead to more severe consequences since forecasters did not trust the predicted precipitation amounts during this extreme scenario and underestimated the flood risk. This gives another motivation why our statistical model and its output in Figure 2 could have been a helpful tool during this flood.
@@
Conditional quantiles (2) are studied in the field of quantile regression, where many flexible methods exist (Meinshausen, 2006; Cannon, 2011; Zhang et al., 2019; Athey et al., 2019). While they work well for quantile levels within the data range, they break down for extreme values of $\tau$ close to 1. Such extreme quantile regression relies on extrapolation results as in (1), where extreme value parameters such as the scale $\sigma(\mathbf{x})$ and shape $\xi(\mathbf{x})$ may be modelled as functions of the covariates through linear models (Wang et al., 2012; Li and Wang, 2019), generalized additive models (Chavez-Demoulin and Davison, 2005; Youngman, 2019) or kernel methods (Daouia et al., 2011; Gardes and Stupfler, 2019; Velthoen et al., 2019). To overcome limitations of additive and kernel-based methods in higher dimensions, more recently, flexible tree-based methods have been combined with the GPD extrapolation for predicting extreme conditional quantiles (Velthoen et al., 2023; Gnecco et al., 2022) or predictive tail distributions (Koh, 2023) on complex data. Tree-based methods have the advantage of requiring little tuning for good prediction performance. However, they can not incorporate additional structure of the data as encountered in time series or spatial applications.
@@
The goal of our work is to combine ideas from extreme value theory and machine learning to propose an extreme quantile regression model that has the ability to extrapolate in the direction of the response $Y$ and to model complex covariate dependencies in the predictors $\mathbf{X}$. We propose an extreme quantile regression network (EQRN) that models covariate-dependent GPD parameters $\sigma(\mathbf{x})$ and $\xi(\mathbf{x})$ as outputs of a neural network. Conditional quantile estimates at the desired extreme level are then readily derived from the estimated conditional tail distribution. Neural networks are known for their ability to model complex dependencies and to approximate any measurable function arbitrarily well (Hornik, 1991). The second advantage is versatility. The deep learning literature is rich in network architectures, activation functions and regularization methods. In particular, convolutions produce shift-invariant models for covariates with spatial dependencies, and recurrent architectures provide models for sequentially dependent observations such as time series. In our application of flood forecasting with time-dependent data, recurrent neural networks (Werbos, 1988; Elman, 1990) are of particular interest.
@@
The main output of our EQRN model for sequentially dependent data is the one-day-ahead risk forecast, either in terms of conditional quantiles or exceedance probabilities. Thanks to the GPD approximation, the method can extrapolate beyond the range of the data, as illustrated in Figure 2. The strength of our recurrent model lies in the ability to exploit information from multiple covariates and capture complex time dependence. It is therefore an effective early warning tool even for unprecedented, record-shattering events, like the 2005 floods. This is of particular importance in a non-stationary system, where climate change makes extreme events increasingly likely (Fischer et al., 2021).
@@
The main contributions of the paper are threefold. Firstly, our EQRN method is the first to model the conditional GPD parameters through neural networks. Thanks to the large number of neural network architectures, this expands the range of possible applications for extreme quantile regression to new areas. While we concentrate on sequentially dependent data, our method can be used, for instance, in combination with convolutional or graphical neural networks for spatial covariates. Secondly, a major technical contribution is to make neural networks applicable in the extreme value context. To stabilize the prediction of extreme quantiles in these highly flexible regression methods, we propose to use an orthogonal reparametrization of the GPD deviance, a suitable choice of activation functions, and the use of the intermediate quantiles as additional covariates; the latter is a new idea that seems to also improve other extreme quantile methods. Finally, a main novelty is the application to flood forecasting and the notion of conditional return levels that can be used as early warnings. With this perspective, our method enables applications in many other areas, such as the one-day-ahead forecasting of the value-at-risk or of the expected shortfall in financial time series.
@@
The paper is organized as follows. In Section 2, we provide background on quantile regression, extreme value theory and neural networks. We propose our EQRN model for both independent and sequentially dependent data in Section 3. Section 4 contains a simulation study to assess the performance of our approach in comparison to existing methods. In Section 5, we describe the Swiss river data, apply our methodology to forecast flood risk and discuss the implications of the results. Section 6 concludes with a brief discussion.
@@ 2.1 Quantile Regression
In the classical quantile regression setup, we observe an independent and identically distributed sample $\mathcal{D}=\left\{(\mathbf{x}_{i},y_{i})\right\}_{i=1}^{n}$ of the random vector $(\mathbf{X},Y)$, where $Y$ is the real-valued response variable and $\mathbf{X}$ is a vector of $p$ covariates (or predictors). One aims at predicting the conditional quantile $Q_{\mathbf{x}}(\tau)$ defined in (2) of $Y$ given $\mathbf{X}=\mathbf{x}$, for some predictor value $\mathbf{x}\in\mathbb{R}^{p}$ and probability level $\tau\in(0,1)$ of interest.
@@
where $\rho_{\tau}(t):=t(\tau-\mathds{1}_{\{t<0\}})$ is the quantile check function (Koenker and Bassett, 1978). Many parametric and non-parametric quantile regression models exist, including linear models (Chernozhukov, 2005), random forests (Athey et al., 2019) and neural networks (Cannon, 2011; Zhang et al., 2019). They yield a conditional quantile estimate by minimizing the empirical quantile loss over the training sample $\mathcal{D}$, that is,
@@
Classical methods for quantile regression that rely on the quantile loss (3) perform well for “moderate” probability levels $\tau$. To define what that means exactly, we typically let $\tau_{n}$ depend on the sample size $n$. The expected number of exceedances of $y_{i}$ over the respective conditional quantile $Q_{\mathbf{x}_{i}}(\tau_{n})$, $i=1,\dots,n$, is given by $n(1-\tau_{n})$. With moderately extreme, or intermediate, we refer to a sequence $\tau_{n}\to 1$ with $n(1-\tau_{n})\to\infty$, meaning that the quantile goes to the upper endpoint of the distribution, but there are more and more exceedances with growing sample size $n$. On the other hand, we call a quantile level $\tau_{n}\to 1$ extreme if $n(1-\tau_{n})\to c\in[0,\infty)$, that is, there are finitely many, or possibly zero, exceedances over $Q_{\mathbf{x}_{i}}(\tau_{n})$ in the sample. In this situation, classical quantile regression methods do not perform well due to the scarcity of observations in the tail of the response.
@@
In the sequel, we omit the dependence on $n$ and write $\tau$ instead of $\tau_{n}$. Intermediate quantile levels will be denoted by $\tau_{0}$.
@@ 2.2 Generalized Pareto Distribution
where the threshold $u(\mathbf{x})$ is chosen as an intermediate quantile $Q_{\mathbf{x}}(\tau_{0})$ at level $\tau_{0}\in(0,1)$ close to 1, and the shape $\xi(\mathbf{x})\in\mathbb{R}$ and scale $\sigma(\mathbf{x})>0$ depend on the covariates; here we omit the dependence of $\sigma(\mathbf{x})$ on the intermediate level $\tau_{0}$ in the notation. This approximation holds under weak conditions on the tail of $Y\mid\mathbf{X}=\mathbf{x}$; see (Balkema and de Haan, 1974; Pickands III, 1975) for the precise statement. This condition is of univariate nature, and it can therefore be verified even in more complex situations, for instance where $\mathbf{X}$ represents the history of a multivariate time series; see Section 4.1 for details. The shape parameter $\xi(\mathbf{x})$ is important since it encodes the tail heaviness of the response: if it is positive, the response has a heavy-tailed distribution such as Pareto or Student-$t$; if it is zero, the response is light-tailed such as a Gaussian or exponential; if it is negative then the response has a finite upper endpoint.
@@
This shows that an estimate $\hat{Q}_{\mathbf{x}}(\tau)$ of an extreme quantile requires estimates of the intermediate quantile $\hat{Q}_{\mathbf{x}}(\tau_{0})$ and of the conditional GPD parameters $\hat{\xi}(\mathbf{x})$ and $\hat{\sigma}(\mathbf{x})$ as functions of the predictor vector. For the intermediate quantile function, we can use any of the existing methods for quantile regression since they work well for this moderate quantile level as discussed above. Estimation of the GPD parameters can be done by specifying a parametric or non-parametric model. We will use neural networks for this purpose, which are introduced in the next section.
The green line in Figure 3 shows estimates $\hat{Q}(\tau)$ for different quantile levels $\tau$ using the approximation (6) without covariate dependence, with empirical intermediate quantile at $\tau_{0}=90\%$ and GPD parameters estimated with maximum likelihood. It can be seen that the extrapolation solves the bias issue of empirical methods.
@@
Figure 3: Left: true $\tau$-quantiles (solid line) compared to empirical estimates (dashed line) and GPD based estimates (dotted line) for moderate to extreme probability levels (log-scale) for sample size 1,000. Estimates are averages over 100 trials. Right: multi-layer perceptron flowchart from input $\mathbf{x}$ to output $g_{\mathcal{W}}(\mathbf{x})$, with loss function $\ell$ and corresponding response $y$.
@@ 2.3 Neural Networks and Conditional Density Estimation
A multi-layer perceptron (MLP) or fully-connected feed-forward neural network model is a parametric family of non-linear functions $g_{\mathcal{W}}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{q}$ that map a $p$-dimensional input $\mathbf{x}$ to a $q$-dimensional output by
@@
where $\mathbf{x}^{(0)}=\mathbf{x}$. The number of hidden layers $L\in\mathbb{N}$, the hidden layer dimensions $h_{1},\ldots,h_{L}\in\mathbb{N}$ and the choice of activation functions $\sigma^{l}:\mathbb{R}^{h_{l}}\rightarrow\mathbb{R}^{h_{l}},l=1,\ldots,L+1$ (with $h_{0}=p$ and $h_{L+1}=q$) are hyperparameters that need to be chosen for instance by cross-validation. The set of trainable parameters to be inferred from data contains all weights and bias terms of the network, that is, $\mathcal{W}=\left\{(W^{l},b^{l});l=1,\ldots,L+1\right\}$, with $W^{l}\in\mathbb{R}^{h_{l}\times h_{l-1}}$ and $b^{l}\in\mathbb{R}^{h_{l}}$. Figure 3 shows a schematic illustration of the transformations inside the MLP.
@@
For conditional density estimation, or distribution regression, we suppose that $Y$ follows a distribution with parametric probability density $f_{Y}(\cdot;{\theta})$ and parameter ${\theta}={\theta}(\mathbf{x})$ depending on the vector $\mathbf{X}=\mathbf{x}$. Conditional density estimation networks are neural networks that aim at outputting conditional estimates $g_{\mathcal{W}}(\mathbf{x})={\theta}(\mathbf{x})$ based on realizations of $\mathbf{X}=\mathbf{x}$ as input (e.g., Cannon, 2012). In this setting, $p$ is the dimension of $\mathbf{X}$ and $q$ is the dimension of ${\theta}$. The loss function is the deviance or negative log-likelihood loss $\ell(y,{{\theta}}(\mathbf{x}))=-\log f_{Y}(y;{{\theta}}(\mathbf{x}))$.
To train a neural network on the training dataset $\mathcal{D}=\left\{(\mathbf{x}_{i},y_{i})\right\}_{i=1}^{n}$ we find the optimal parameter values minimizing the average empirical loss, that is,
@@
This is generally achieved via backpropagation using mini-batches and optimization algorithms such as the well-performing gradient descent variants (Kingma and Ba, 2014; Tieleman and Hinton, 2012; Duchi et al., 2011). Since neural networks are typically overparameterized, overfitting has to be prevented with regularization methods such as $L_{2}$ weight penalties for narrow networks and dropout (Srivastava et al., 2014) for deeper architectures. As the optimization problem (8) is often non-convex, local-minima convergence is an issue. Restricting training to a subset of $\mathcal{D}$ and keeping the rest to track the validation loss at the end of each epoch helps to avoid local minima by learning rate decay. Restarting training with different initializations and keeping the best fit in terms of validation loss often leads to lower minima. Early stopping based on the validation loss is another measure against overfitting. The final validation loss is used for model selection and the choice of optimal hyperparameters; for more details on the fitting of neural networks, see Goodfellow et al. (2016).
When observations are dependent in space or time, generalizations of the MLP exist to account for these particular structures. Convolutional and graph neural networks exploit neighbourhood information with parsimonious architectures that are effective for images, graphs or spatial observations (LeCun et al., 2015; Scarselli et al., 2009). We concentrate here on methods for sequential dependence that arises typically in time series $\left\{(\mathbf{X}_{t},Y_{t})\right\}_{t=1}^{T}$. For this type of data, recurrent architectures of the network allow capturing dependence between observations. A simple recurrent neural network (RNN) layer (Werbos, 1988; Elman, 1990) takes as input a vector $\mathbf{x}_{t}$ and outputs the hidden recurrent state vector
@@
depending both on $\mathbf{x}_{t}$ and the hidden state $\mathbf{h}_{t-1}$ recursively resulting from the previous inputs $\mathbf{x}_{t-1}$ and $\mathbf{h}_{t-2}$ in the sequence. Here and in the sequel, the bias vectors $b_{\cdot}$ and weight matrices $W_{\cdot\cdot}$, indexed by the input and output variables, are the trainable parameters. This model has then been improved by the addition of a gating cell state $\mathbf{c}_{t}$, to avoid vanishing gradient issues, as well as a forget gate $\mathbf{f}_{t}$, an input gate $\mathbf{i}_{t}$ and output gate $\mathbf{o}_{t}$, to control both short- and long-term dependencies in the sequence. This yields the long short-term memory (LSTM) layer (Hochreiter and Schmidhuber, 1997; Gers et al., 2000, 2003; Jozefowicz et al., 2015)
@@
where $\sigma(z)=1/(1+\exp(-z))$ is the sigmoid activation and $\odot$ is the Hadamard (or componentwise) product; see Figure 4 for an illustration. The common dimension of the vectors defined in (9) is a hyperparameter of the layer. The input of this layer $\tilde{\mathbf{x}}_{t}:=(\mathbf{x}_{t-s},\ldots,\mathbf{x}_{t-1})$ can include predictors from the past to model longer dependencies, where $s\in\mathbb{N}$ determines the time horizon.
The LSTM model has been simplified by Cho et al. (2014) into the gated recurrent unit (GRU) layer, which has become a popular alternative. A multi-layer recurrent network is obtained by considering the $\mathbf{h}_{t}$ as a sequence of inputs for the following recurrent layers; see Figure S.1 in Supplementary Material S.1. Usually, a fully connected layer as in (7) is used to map the hidden state of the final recurrent layer to the network output $\tilde{g}_{\mathcal{W}}(\tilde{\mathbf{x}}_{t})$.

### v4

@@ 1 Introduction
Figure 1 shows the river catchment of a gauging station on the Aare river in Bern, Switzerland. It is part of the Aare–Rhein basin, where flooding is a major economic and safety concern (Andres et al., 2021). The Swiss Federal Office for the Environment (FOEN) provides recordings of daily average discharges throughout the country. For risk assessment they report the 100-year return levels using the GEV method for annual maxima and the GPD for large daily discharges. The horizontal lines in Figure 2 show estimates of this return level at the Bernese station based on data from the years 1930–1958. Both methods give very similar results.
@@
The disadvantage of such an unconditional approach is twofold. First, the return level is static and unable to reflect changes in the size of extreme floods over time, which can occur, for instance, due to climate change. For the Bernese station, for instance, a structural break has been observed in the nineties without a clearly defined cause [^1] See flood report of the FOEN at https://www.hydrodaten.admin.ch/en/2135.html., making classical extreme value modelling challenging. Second, while the return level $Q^{T}$ is relevant for the construction of long-term flood infrastructure, it can not be used to assess the risk of flooding on a given day. Such forecasting of extreme events is crucial for early warning systems. Indeed, the probability of exceeding on a particular day a given high threshold, say the (constant) 100-year return level, depends on many covariates $\mathbf{X}$ such as the river flows upstream and precipitation in the catchments during the preceding days and weeks.
@@
The top panel of Figure 2 shows one-day-ahead forecasts of such conditional 100-year quantiles $Q^{100}_{\mathbf{x}}$ for the Bernese river data from the method of this paper, fitted to the years 1930–1958. We see that, as opposed to the unconditional return level $Q^{100}$, the conditional quantile changes from day to day, depending on past precipitation and river flows. In fact, on August 21, the day before the first exceedance of the unconditional $Q^{100}$ during the serious flood of August 2005 in Bern, the size of the conditional 100-year event predicted for the next day was much higher than on other days, and the forecasted conditional probability of such an exceedance (bottom panel of Figure 2) was 920 times larger than the static 100-year probability. Both outputs can be used as triggers for early warnings and additional flood management measures. In particular, the days when an exceedance above $Q^{100}$ is likely can be effectively pinpointed thanks to their temporal sparsity in the probability forecast.
@@
Forecasting extreme events is notoriously difficult due to the low occurrence probabilities involved. This is particularly true for hydrological models, which are typically used by national agencies and which do not use explicit tail extrapolation. In the aftermath of the 2005 flood, the FOEN published a detailed analysis of the internal forecasting procedures during this event (Bezzola and Hegg, 2007). The report showed that too-late warnings lead to more severe consequences since forecasters did not trust the predicted precipitation amounts during this extreme scenario and underestimated the flood risk. This gives another motivation why our statistical model and its output in Figure 2 could have been a helpful tool during this flood.
@@
Conditional quantiles (2) are studied in the field of quantile regression, where many flexible methods exist (Meinshausen, 2006; Cannon, 2011; Zhang et al., 2019; Athey et al., 2019). While they work well for quantile levels within the data range, they break down for extreme values of $\tau$ close to 1. Such extreme quantile regression relies on extrapolation results as in (1), where extreme value parameters, such as the scale $\sigma(\mathbf{x})$ and shape $\xi(\mathbf{x})$, may be modelled as functions of the covariates through linear models (Wang et al., 2012; Li and Wang, 2019), generalized additive models (Chavez-Demoulin and Davison, 2005; Youngman, 2019), or kernel methods (Daouia et al., 2011; Gardes and Stupfler, 2019; Velthoen et al., 2019). To overcome limitations of additive and kernel-based methods in higher dimensions, more recently, flexible tree-based methods have been combined with the GPD extrapolation for predicting extreme conditional quantiles (Velthoen et al., 2023; Gnecco et al., 2022) or predictive tail distributions (Koh, 2023) on complex data. Tree-based methods have the advantage of requiring little tuning for good prediction performance. However, they can not incorporate additional structure of the data as encountered in time series or spatial applications.
@@
The goal of our work is to combine ideas from extreme value theory and machine learning to propose an extreme quantile regression model that has the ability to extrapolate in the direction of the response $Y$ and to model complex covariate dependencies in the predictors $\mathbf{X}$. We propose an extreme quantile regression network (EQRN) that models covariate-dependent GPD parameters $\sigma(\mathbf{x})$ and $\xi(\mathbf{x})$ as outputs of a neural network. Conditional quantile estimates at the desired extreme level are then readily derived from the estimated conditional tail distribution. Neural networks are known for their ability to model complex dependencies and to approximate any measurable function arbitrarily well (Hornik, 1991). The second advantage is versatility. The deep learning literature is rich in network architectures, activation functions, and regularization methods. In particular, convolutions produce shift-invariant models for covariates with spatial dependencies, and recurrent architectures provide models for sequentially dependent observations such as time series. In our application of flood forecasting with time-dependent data, recurrent neural networks (Werbos, 1988; Elman, 1990) are of particular interest.
@@
The main output of our EQRN model for sequentially dependent data is the one-day-ahead risk forecast, either in terms of conditional quantiles or exceedance probabilities. Thanks to the GPD approximation, the method can extrapolate beyond the range of the data, as illustrated in Figure 2. The strength of our recurrent model lies in the ability to exploit information from multiple covariates and capture complex time dependence. It is, therefore, an effective early warning tool even for unprecedented, record-shattering events, like the 2005 floods. This is of particular importance in a nonstationary system, where climate change makes extreme events increasingly likely (Fischer et al., 2021).
@@
The main contributions of the paper are threefold. First, our EQRN method is the first to model the conditional GPD parameters through neural networks. Thanks to the large number of neural network architectures, this expands the range of possible applications for extreme quantile regression to new areas. While we concentrate on sequentially dependent data, our method can be used, for instance, in combination with convolutional or graphical neural networks for spatial covariates. Second, a major technical contribution is to make neural networks applicable in the extreme value context. To stabilize the prediction of extreme quantiles in these highly flexible regression methods, we propose to use an orthogonal reparametrization of the GPD deviance, a suitable choice of activation functions, and the use of the intermediate quantiles as additional covariates; the latter is a new idea that seems to also improve other extreme quantile methods. Finally, a main novelty is the application to flood forecasting and the notion of conditional return levels that can be used as early warnings. With this perspective our method enables applications in many other areas, such as the one-day-ahead forecasting of the value-at-risk or of the expected shortfall in financial time series.
@@
The paper is organized as follows. In Section 2 we provide background on quantile regression, extreme value theory, and neural networks. We propose our EQRN model for both independent and sequentially dependent data in Section 3. Section 4 contains a simulation study to assess the performance of our approach in comparison to existing methods. In Section 5 we describe the Swiss river data, apply our methodology to forecast flood risk and discuss the implications of the results. Section 6 concludes with a brief discussion.
@@ 2.1 Quantile regression
In the classical quantile regression setup, we observe an independent and identically distributed sample $\mathcal{D}=\left\{(\mathbf{x}_{i},y_{i})\right\}_{i=1}^{n}$ of the random vector $(\mathbf{X},Y)$, where $Y$ is the real-valued response variable and $\mathbf{X}$ is a vector of $p$ covariates (or predictors). One aims at predicting the conditional quantile $Q_{\mathbf{x}}(\tau)$ defined in (2) of $Y$, given $\mathbf{X}=\mathbf{x}$, for some predictor value $\mathbf{x}\in\mathbb{R}^{p}$ and probability level $\tau\in(0,1)$ of interest.
@@
where $\rho_{\tau}(t):=t(\tau-\mathds{1}_{\{t<0\}})$ is the quantile check function (Koenker and Bassett, 1978). Many parametric and nonparametric quantile regression models exist, including linear models (Chernozhukov, 2005), random forests (Athey et al., 2019), and neural networks (Cannon, 2011; Zhang et al., 2019). They yield a conditional quantile estimate by minimizing the empirical quantile loss over the training sample $\mathcal{D}$, that is,
@@
Classical methods for quantile regression that rely on the quantile loss (3) perform well for “moderate” probability levels $\tau$. To define what that means exactly, we typically let $\tau_{n}$ depend on the sample size $n$. The expected number of exceedances of $y_{i}$ over the respective conditional quantile $Q_{\mathbf{x}_{i}}(\tau_{n})$, $i=1,\dots,n$, is given by $n(1-\tau_{n})$. With moderately extreme, or intermediate, we refer to a sequence $\tau_{n}\to 1$ with $n(1-\tau_{n})\to\infty$, meaning that the quantile goes to the upper endpoint of the distribution, but there are more and more exceedances with growing sample size $n$. On the other hand, we call a quantile level $\tau_{n}\to 1$ extreme if $n(1-\tau_{n})\to c\in[0,\infty)$; that is, there are finitely many, or possibly zero, exceedances over $Q_{\mathbf{x}_{i}}(\tau_{n})$ in the sample. In this situation, classical quantile regression methods do not perform well due to the scarcity of observations in the tail of the response.
@@
In the sequel we omit the dependence on $n$ and write $\tau$ instead of $\tau_{n}$. Intermediate quantile levels will be denoted by $\tau_{0}$.
@@ 2.2 Generalized Pareto distribution
where the threshold $u(\mathbf{x})$ is chosen as an intermediate quantile $Q_{\mathbf{x}}(\tau_{0})$ at level $\tau_{0}\in(0,1)$ close to 1, and the shape $\xi(\mathbf{x})\in\mathbb{R}$ and scale $\sigma(\mathbf{x})>0$ depend on the covariates; here we omit the dependence of $\sigma(\mathbf{x})$ on the intermediate level $\tau_{0}$ in the notation. This approximation holds under weak conditions on the tail of $Y\mid\mathbf{X}=\mathbf{x}$; see Balkema and de Haan (1974); Pickands III (1975) for the precise statement. This condition is of univariate nature, and it can, therefore, be verified even in more complex situations, for instance, where $\mathbf{X}$ represents the history of a multivariate time series; see Section 4.1 for details. The shape parameter $\xi(\mathbf{x})$ is important since it encodes the tail heaviness of the response: if it is positive, the response has a heavy-tailed distribution such as Pareto or Student-$t$; if it is zero, the response is light-tailed such as a Gaussian or exponential; if it is negative, then the response has a finite upper endpoint.
@@
This shows that an estimate $\hat{Q}_{\mathbf{x}}(\tau)$ of an extreme quantile requires estimates of the intermediate quantile $\hat{Q}_{\mathbf{x}}(\tau_{0})$ and of the conditional GPD parameters $\hat{\xi}(\mathbf{x})$ and $\hat{\sigma}(\mathbf{x})$ as functions of the predictor vector. For the intermediate quantile function, we can use any of the existing methods for quantile regression since they work well for this moderate quantile level, as discussed above. Estimation of the GPD parameters can be done by specifying a parametric or nonparametric model. We will use neural networks for this purpose, which are introduced in the next section.
The green line in Figure 3 shows estimates $\hat{Q}(\tau)$ for different quantile levels $\tau$, using the approximation (6) without covariate dependence, with empirical intermediate quantile at $\tau_{0}=90\%$ and GPD parameters estimated with maximum likelihood. It can be seen that the extrapolation solves the bias issue of empirical methods.
@@
Figure 3: Left: True $\tau$-quantiles (solid line) compared to empirical estimates (dashed line) and GPD based estimates (dotted line) for moderate to extreme probability levels (log-scale) for sample size 1,000. Estimates are averages over 100 trials. Right: Multilayer perceptron flowchart from input $\mathbf{x}$ to output $g_{\mathcal{W}}(\mathbf{x})$, with loss function $\ell$ and corresponding response $y$.
@@ 2.3 Neural networks and conditional density estimation
A multilayer perceptron (MLP) or fully-connected feed-forward neural network model is a parametric family of nonlinear functions $g_{\mathcal{W}}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{q}$ that map a $p$-dimensional input $\mathbf{x}$ to a $q$-dimensional output by
@@
where $\mathbf{x}^{(0)}=\mathbf{x}$. The number of hidden layers $L\in\mathbb{N}$, the hidden layer dimensions $h_{1},\ldots,h_{L}\in\mathbb{N}$, and the choice of activation functions $\sigma^{l}:\mathbb{R}^{h_{l}}\rightarrow\mathbb{R}^{h_{l}},l=1,\ldots,L+1$ (with $h_{0}=p$ and $h_{L+1}=q$) are hyperparameters that need to be chosen, for instance, by cross-validation. The set of trainable parameters to be inferred from data contains all weights and bias terms of the network, that is, $\mathcal{W}=\left\{(W^{l},b^{l});l=1,\ldots,L+1\right\}$, with $W^{l}\in\mathbb{R}^{h_{l}\times h_{l-1}}$ and $b^{l}\in\mathbb{R}^{h_{l}}$. Figure 3 shows a schematic illustration of the transformations inside the MLP.
@@
For conditional density estimation, or distribution regression, we suppose that $Y$ follows a distribution with parametric probability density $f_{Y}(\cdot;{\theta})$ and parameter ${\theta}={\theta}(\mathbf{x})$ depending on the vector $\mathbf{X}=\mathbf{x}$. Conditional density estimation networks are neural networks that aim at outputting conditional estimates $g_{\mathcal{W}}(\mathbf{x})={\theta}(\mathbf{x})$ based on realizations of $\mathbf{X}=\mathbf{x}$ as input (e.g., Cannon, 2012). In this setting, $p$ is the dimension of $\mathbf{X}$, and $q$ is the dimension of ${\theta}$. The loss function is the deviance or negative log-likelihood loss $\ell(y,{{\theta}}(\mathbf{x}))=-\log f_{Y}(y;{{\theta}}(\mathbf{x}))$.
To train a neural network on the training dataset $\mathcal{D}=\left\{(\mathbf{x}_{i},y_{i})\right\}_{i=1}^{n}$, we find the optimal parameter values minimizing the average empirical loss, that is,
@@
This is generally achieved via backpropagation using mini-batches and optimization algorithms, such as the well-performing gradient descent variants (Kingma and Ba, 2014; Tieleman and Hinton, 2012; Duchi et al., 2011). Since neural networks are typically overparameterized, overfitting has to be prevented with regularization methods, such as $L_{2}$ weight penalties for narrow networks and dropout (Srivastava et al., 2014) for deeper architectures. As the optimization problem (8) is often nonconvex, local-minima convergence is an issue. Restricting training to a subset of $\mathcal{D}$ and keeping the rest to track the validation loss at the end of each epoch helps to avoid local minima by learning rate decay. Restarting training with different initializations and keeping the best fit in terms of validation loss often leads to lower minima. Early stopping based on the validation loss is another measure against overfitting. The final validation loss is used for model selection and the choice of optimal hyperparameters; for more details on the fitting of neural networks, see Goodfellow et al. (2016).
When observations are dependent in space or time, generalizations of the MLP exist to account for these particular structures. Convolutional and graph neural networks exploit neighbourhood information with parsimonious architectures that are effective for images, graphs, or spatial observations (LeCun et al., 2015; Scarselli et al., 2009). We concentrate here on methods for sequential dependence that arises typically in time series $\left\{(\mathbf{X}_{t},Y_{t})\right\}_{t=1}^{T}$. For this type of data, recurrent architectures of the network allow capturing dependence between observations. A simple recurrent neural network (RNN) layer (Werbos, 1988; Elman, 1990) takes as input a vector $\mathbf{x}_{t}$ and outputs the hidden recurrent state vector
@@
depending both on $\mathbf{x}_{t}$ and the hidden state $\mathbf{h}_{t-1}$ recursively resulting from the previous inputs $\mathbf{x}_{t-1}$ and $\mathbf{h}_{t-2}$ in the sequence. Here and in the sequel, the bias vectors $b_{\cdot}$ and weight matrices $W_{\cdot\cdot}$, indexed by the input and output variables, are the trainable parameters. This model has then been improved by the addition of a gating cell state $\mathbf{c}_{t}$, to avoid vanishing gradient issues, as well as a forget gate $\mathbf{f}_{t}$, an input gate $\mathbf{i}_{t}$, and output gate $\mathbf{o}_{t}$, to control both short- and long-term dependencies in the sequence. This yields the long short-term memory (LSTM) layer (Hochreiter and Schmidhuber, 1997; Gers et al., 2000, 2003; Jozefowicz et al., 2015)
@@
where $\sigma(z)=1/(1+\exp(-z))$ is the sigmoid activation and $\odot$ is the Hadamard (or componentwise) product; see Figure 4 for an illustration. The common dimension of the vectors defined in (9), is a hyperparameter of the layer. The input of this layer $\tilde{\mathbf{x}}_{t}:=(\mathbf{x}_{t-s},\ldots,\mathbf{x}_{t-1})$ can include predictors from the past to model longer dependencies, where $s\in\mathbb{N}$ determines the time horizon.
The LSTM model has been simplified by Cho et al. (2014) into the gated recurrent unit (GRU) layer, which has become a popular alternative. A multilayer recurrent network is obtained by considering the $\mathbf{h}_{t}$ as a sequence of inputs for the following recurrent layers; see Figure S.1 in Supplementary Material S.1. Usually, a fully connected layer as in (7) is used to map the hidden state of the final recurrent layer to the network output $\tilde{g}_{\mathcal{W}}(\tilde{\mathbf{x}}_{t})$.